# How to be an orthodox quantum mechanic

### v1

@@ 5 Results and discussion
Figure 1: Mode frequency of the responses to each question. The bars represent different samples: all (black), undergraduate (yellow), graduate (blue), Wikipedia notable (red), and modern (grey).
Table 3: Degree of consensus from the full sample. $p$ is the probability of obtaining the given degree of consensus by chance from a random, binomial distribution.
Assertion Affirmation $p$ $\sigma$
Collapse 0.77 1.62e-04 3.1
Non-ensemble psi 0.95 1.16e-10 5.2
Ontologically complete 0.85 5.46e-06 4.0
No trajectories 0.45 3.64e-01 0.5
Complementarity 0.53 2.43e-01 0.3
Wave-particle duality 0.79 1.62e-04 3.3
Single particle uncertainty 0.82 3.31e-05 3.7
Single particle wave-like 0.85 5.46e-06 4.0
Measurement creates values 0.71 2.28e-03 2.4
Probability from uncertainty 0.35 8.14e-02 1.7
Measurement disturbs system 0.92 6.54e-08 4.9
Uncertainty linked to disturbance 0.71 2.28e-03 2.4
Table 4: Degree of consensus from the undergraduate sample. $p$ is the probability of obtaining the given degree of consensus by chance from a random, binomial distribution.
Assertion Affirmation $p$ $\sigma$
Collapse 0.76 9.61e-03 2.3
Non-ensemble psi 0.97 1.91e-06 4.1
Ontologically complete 0.79 2.21e-03 2.5
No trajectories 0.39 3.24e-01 0.9
Complementarity 0.55 1.80e-01 0.5
Wave-particle duality 0.84 3.64e-04 3.0
Single particle uncertainty 0.82 3.64e-04 2.8
Single particle wave-like 0.89 3.81e-05 3.4
Measurement creates values 0.76 9.61e-03 2.3
Probability from uncertainty 0.39 3.24e-01 0.9
Measurement disturbs system 0.97 1.91e-06 4.1
Uncertainty linked to disturbance 0.71 9.61e-03 1.8
Table 5: Degree of consensus from the graduate sample. $p$ is the probability of obtaining the given degree of consensus by chance from a random, binomial distribution.
Assertion Affirmation $p$ $\sigma$
Collapse 0.79 6.47e-03 2.1
Non-ensemble psi 0.93 6.10e-05 3.2
Ontologically complete 0.93 6.10e-05 3.2
No trajectories 0.54 2.12e-01 0.3
Complementarity 0.50 3.95e-01 0.0
Wave-particle duality 0.71 2.87e-02 1.6
Single particle uncertainty 0.82 9.16e-04 2.4
Single particle wave-like 0.79 6.47e-03 2.1
Measurement creates values 0.64 8.98e-02 1.1
Probability from uncertainty 0.29 8.98e-02 1.6
Measurement disturbs system 0.86 9.16e-04 2.7
Uncertainty linked to disturbance 0.71 2.87e-02 1.6
Table 6: Degree of consensus from the ‘Wikipedia notable’ sample. $p$ is the probability of obtaining the given degree of consensus by chance from a random, binomial distribution.
Assertion Affirmation $p$ $\sigma$
Collapse 0.74 2.45e-02 1.9
Non-ensemble psi 0.94 7.63e-06 3.6
Ontologically complete 0.88 1.37e-04 3.2
No trajectories 0.35 1.66e-01 1.2
Complementarity 0.41 3.15e-01 0.7
Wave-particle duality 0.71 2.45e-02 1.7
Single particle uncertainty 0.82 1.17e-03 2.7
Single particle wave-like 0.76 6.36e-03 2.2
Measurement creates values 0.74 2.45e-02 1.9
Probability from uncertainty 0.32 1.66e-01 1.5
Measurement disturbs system 0.94 7.63e-06 3.6
Uncertainty linked to disturbance 0.79 1.17e-03 2.4
Table 7: Degree of consensus from the ‘modern’ sample. $p$ is the probability of obtaining the given degree of consensus by chance from a random, binomial distribution.
Assertion Affirmation $p$ $\sigma$
Collapse 0.71 1.13e-02 2.0
Non-ensemble psi 0.90 1.49e-06 3.9
Ontologically complete 0.79 7.72e-04 2.9
No trajectories 0.35 7.58e-02 1.4
Complementarity 0.40 2.71e-01 1.0
Wave-particle duality 0.71 1.13e-02 2.0
Single particle uncertainty 0.60 1.54e-01 1.0
Single particle wave-like 0.75 3.31e-03 2.4
Measurement creates values 0.67 3.20e-02 1.6
Probability from uncertainty 0.23 1.13e-02 2.7
Measurement disturbs system 0.85 1.39e-04 3.5
Uncertainty linked to disturbance 0.48 5.81e-01 0.2
We begin with propositions related to the nature of the quantum state. Interestingly, we see that not all the books regard $\psi$ as ontologically complete. However, the support for this proposition is robust with assent frequency $0.85$ ($p=5.46\times 10^{-6}$) in the full sample, with all of sub-samples having assent frequency $\geq 0.79$ ($p\leq 7.72\times 10^{-4}$). It is interesting that the lowest support for this occurs within the undergraduate and ‘modern’ samples. In terms of dissent, Phillips [24] and Moore [27] limit themselves to saying that $\psi$ can determine all the probabilities associated with a system but do not go on to say this is all the information allowed. Whereas, [17] openly suggests that interpretive difficulties might be overcome by finding a more general theory to supersede quantum mechanics. We also see strong support (a minimum frequency of $0.9$ with $p\leq 6.1\times 10^{-5}$) for the notion that a quantum state describes a single instance (as opposed to an ensemble). Notably, this achieves $0.95$ ($p=1.16\times 10^{-10}$) assent in the full sample and the undergraduate works are near unanimous on this count. Support is weaker for state collapse, where we obtain sub-sample assent frequencies above $0.71$ ($p\leq 2.45\times 10^{-2}$) while achieving $f=0.77$ ($p=1.62\times 10^{-4}$) in the full sample. From these results we can also see that an orthodox view requires the state to contain the entire ontological situation and apply to individual particles. This, in turn, requires us to believe that superposition states can be occupied by individual systems. Of the three propositions, only affirmation of collapse reduces slightly if we consider sources after 1980. Thus, the orthodoxy on these issues seems largely unchanged over the last century. It is noteworthy that, in a sample of $33$ works, there are only two dissenters to the notion that quantum states collapse: Weinberg [17] as well as Binney & Skinner [18]. Additionally, the topic is not mentioned by Griffiths, Peres, Müller-Kirsten, and Schiff [28, 29, 22, 46]. Similarly, only Peres [29] dissents from the notion that quantum states describe individual instances. Interestingly, we note that Home [4] argues that collapse is not essential to characterise the orthodoxy. However, this is strongly disputed by the data gathered here. Overall, we see that the orthodoxy has it that the quantum state is ontologically complete, refers to individual instances of systems, and collapses on measurement.
Now we turn to propositions on the nature of particles. A strong degree of consensus is clear around the notion that individual particles display wave-like behaviour with assent frequency $0.85$ at ($p\leq 5.46\times 10^{-6}$), sub-sample frequency $\geq 0.75$ ($p\leq 6.36\times 10^{-3}$). The only dissenters here being Weinberg [17], Binney & Skinner [18], and Sakurai [33]. The topic is also not discussed in von Neumann [44] (perhaps unsurprising given the mathematical focus) and Griffiths [28].
Reasonable support also exists for the notion of wave-particle duality with assent $0.79$ ($p\leq 1.62\times 10^{-4}$) in the full sample, sub-sample frequency $\geq 0.71$ ($p\leq 2.87\times 10^{-2}$). Here only von Neumann [44] and Binney & Skinner [18] outright denying this proposition. It is also not mentioned by Weinberg [17], Mandl [30], Townsend [31], and Sakurai [33].
For the proposition that individual particles being subject to the uncertainty principle, that is, one cannot know a given particle’s position and momentum (for example) simultaneously to arbitrary precision, we see more controversy. This is so because the full sample finds assent of $0.82$ ($p=3.31\times 10^{-5}$), however, the ‘modern’ subs-sample only registers $0.6$ ($p=0.154$). The reason for this change is that Griffiths [28], Peres [29], McIntyre [20], Moore [27], and Sakurai [33] all actively deny this proposition. This shows some evidence of a change in the consensus over time. However, a new orthodoxy is not yet entrenched, as Binney & Skinner [18], Longair [19], Townsend [31], Konishi & Paffuti [21], Omnés [26], Shankar [34], Phillips [24], Greiner [32], and Le Bellac [23] all affirm the proposition. Additionally, it is reasonably inferred that Mandl [30] and Weinberg [17] also support it. Home [4] does not include this in their conception of the orthodoxy on the grounds that is a flawed proposition. As argued in [7], the uncertainty principle could not function in the suggested manner. If the measurements from individual particles had the uncertainty principle’s scatter, the final ensemble distribution would have a larger uncertainty. Thus, preventing the verification of the principle as derived, where it applies to standard deviations over an ensemble of trials. Despite scepticism on this point, it is clearly wrong to exclude this from the orthodoxy, as the idea is uniformly affirmed by older works and still prevalent, but less so, in modern texts.
It is notable that affirmation of complementarity is generally weak, even the full sample only finds an agreement frequency of $0.53$ ($p=0.243$). This is not due to denial but simply it never being mentioned in many later texts. Regardless, it cannot have a place in any robust orthodoxy. We also see that the issue of whether particles possess trajectories is contentious, with frequency $0.45$ ($p=0.364$) in the full sample. This becomes even more contentious in the ‘modern’ sample with $0.35$ ($p=0.0758$). This proposition still finds agreement in modern works like Konishi & Paffuti [21], Cohen-Tannoudji [36], Shankar [34], Greiner [32], and Phillips [24], with partial agreement in Moore [27] and Longair [19]. However, the mode response is 00 in the undergraduate, ‘Wikipedia notable’ and ‘modern’ samples. So there is some indication that this proposition is increasingly marginalised by the orthodoxy. It might also follow that those who affirm the “single particle uncertainty” proposition should, for consistency, agree that particles have no trajectory (at least for resolutions finer than that allowed by uncertainty). However, we have tried to avoid forcing logical consistency on authors where there is some ambiguity or nuance.
@@
Next, we turn to propositions around the question of why quantum mechanics predicts only statistical results. There is clear consensus, regardless of whether or not we confine our attention to modern works, that measurement disturbs quantum systems with assent $0.92$ ($p=6.54\times 10^{-8}$) in the full sample. Notably, the undergraduate sample has a frequency of $\sim 1.0$ for this proposition and only the graduate and ‘modern’ samples drop below $0.94$. Indeed, only Peres [29] actively dissents. Their dissent is justified by the possibility of non-interactive measurement in QM [48]. There is also support for the idea that measurement itself actually creates the observed values with frequency $0.71$ ($p=2.28\times 10^{-3}$) in the full sample while retaining $0.67$ ($p\leq 3.2\times 10^{-2}$) in the ‘modern’ sub-sample. It should be noted that the affirmation is only stated openly 16 out of 29 times, despite being a plausible corollary to “measurement disturbs quantum systems” and that there are no active dissenters, not even [29]. The low significance thus represents the number of times the assent was only inferred. Only three authors do not provide grounds for inferring agreement: Sakurai [33], Mandl [30], and Weinberg [17]. The common case for inference is as follows: $\psi$ applies to individual systems, hence superpositions are real. Measurement disturbs the system and changes a superposition to an eigenstate, so it is either changing an unknown underlying value or creating a definite one where there was none. In either case the pre-existing value is totally superfluous if the collapse process is irreversible and indeterminist.
The notion that probabilities in QM result from the uncertainty principle has always been a minority view it appears, as most authors do not mention anything like it (this question has a mode of 00 in all samples). There is still some support for the notion that the uncertainty principle is a consequence of measurement disturbance, with assent frequency $>0.71$ outside the ‘modern’ sample where it is only $0.48$. This indicates that affirmation of this proposition may be falling out of the orthodoxy. Interestingly, the sample with most affirmation for this last proposition is the “Wikipedia notable” one, with $0.79$ assent ($p=1.17\times 10^{-3}$).
@@
Thus, for the orthodoxy, we do not reveal the values of studied quantum systems but force them to assume definite values (or alter them) via some uncontrolled disturbance attached to the abstract notion of measurement. The “uncontrolled” aspect of this is vital to square the statistical nature of QM with the proposition that quantum states apply to individual instances of systems. Here we see emergence of anti-realism as a pillar of the orthodoxy. This is entailed by the fact that we are not exploring the pre-existing properties of quantum systems so much as forcing them into altered states under measurement.
To better understand the development of the orthodoxy over time we can also consult the mean responses in each sub-sample, as in Figure 2. Here we can see the modern sample has a consistently lower mean response to the vast majority of propositions. The most significant difference occurs for the “single particle uncertainty” proposition, with a $40$% reduction in the mean value relative to the other samples. This reduction is twice the difference in assent frequency between the samples, indicating the presence of active dissent from the proposition. This cannot be seen from the assent frequency, as $-1$ and 00 have the same effect on the frequency of $1$. The proposition of “ontological completeness” sees a decrease in mean response of around 20% in the ‘modern’ sub-sample that is not evident in response frequency, again showing evidence of reduced consensus on this point. The “uncertainty from disturbance” proposition also has a mean response decline of $30$%. However, this correlates well with the assent frequency, indicating only that this idea is increasingly not mentioned in more recent works. These demonstrate that there are propositions that were previously orthodox (with assent frequencies above $0.7$ in other samples) that have become more controversial over time. Some less popular positions (“probability from uncertainty” and “no trajectories”) also exhibit evidence of increased controversy in the modern sample, showing smaller mean values than even other samples where the mode is similarly 00.
Figure 2: Average values of the responses to each question. The bars represent different samples: all (black), undergraduate (yellow), graduate (blue), Wikipedia notable (red), and modern (grey).
@@
where the sum runs over the propositions, $w_{i}$ is the average response value, and $a_{i}$ is that of the reference work in question. Note that we exclude propositions that do not have assent frequency $>0.6$.
The average orthodoxity is $74\pm 25$%, indicating a robust degree of agreement on average. For convenience we order Fig. 3 by year of first publication. This allows us to determine if there are any trends in the orthodoxy. What we see is quite remarkable. Early works adhere very closely to the orthodoxy (as might be expected). The period around the early 90’s shows a fluctuation of far lower agreement with the orthodoxy. However, the orthodoxy reasserts itself after this point (but slightly more weakly than in the earlier works). The 90’s dip in the orthodoxity is constituted by the works: Peres [29], Moore [27], Griffiths [28], Mandl [30], Rae [13], and Sakurai [33]. Overall, there is some evidence for an evolutionary trend in the orthodoxy, but it must be admitted that modern works quite closely follow the interpretation put forward by their predecessors. Despite this as an overall statement, we can see recent dissent from Weinberg [17] and some clearly unorthodox positions in both Binney & Skinner [18] and Griffiths [28]. This is of interest given the prominence of the first’s author and that the latter two also appear on the “Wikipedia Notable” list. It is also interesting that the modern reduction in orthodoxity (even if unspectacular on average) revolves around the questions of the wave-like behaviour of individual particles. Something that is denied by Weinberg, Binney & Skinner, Griffiths, and Sakurai [17, 18, 28, 33], these being the works with some of the lowest orthodoxity. The lack of orthodoxity for Peres [29] is rooted in the author’s instrumentalism (being far closer to the version of a Copenhagen viewpoint as understood by Stapp [2]).
@@
For additional works that were not included in the sample we find the following orthodoxity measures: Ballentine [12] $-0.89$, Home [4] 0.44, Hannabuss [14] 0.43, and Nielsen & Chuang [15] 0.44. The last of these, Nielsen & Chuang [15], only shows a lower orthodoxity by virtue of never mentioning the topic of waves. Their strong endorsement of Peres [29] and Sakurai [33] might indicate little sympathy with this aspect of the orthodoxy, but the clarity of this is marred by also endorsing Cohen-Tannoudji [36]. Interestingly, Home’s characterisation of the orthodoxy [4] appears incomplete at $0.44$. Additionally, this is only when one concentrates on the aspects attributed to the non-ensemble case. Home considers ensemble views as part of the orthodoxy on the basis that they view the formalism as being complete (there are no supplemental entities needed to explain what is going on). This principle as the central pillar of the orthodoxy is apriori dubious. This is because the notion of collapse is a supplement to the basic mathematical formalism [49, 12, 17], in the form of an additional physical/quasi-physical process with attendant projection postulate. Since Copenhagen-like interpretations hinge so heavily on collapse, their popularity implies that Home’s primary condition of orthodoxy becomes more doubtful. Interestingly, Ballentine [12], who would express an orthodox view according to Home, achieves an orthodoxity of $-0.89$ here, dissenting from (or not addressing) every proposition considered. This is because Home’s rationale for including ensemble views in the orthodoxy may not be well justified. Home argues that, to enjoy a complete formalism, the ensemble must be supplemented by hidden variables that explain the assignment of definite values to individual realisations. However, an ensemble view only requires statistical completeness, $\psi$ can answer all statistical questions, not true ontological completeness. In this regard Home forces ensemble views to be able to determine the behaviour of single instances, in contradiction with their basic premise. In addition, no such definite assignment is required of other statistical theories in science (which usually exist as a limit of some other, more detailed theory). We do not require statistical mechanics to explain how velocities are assigned to the molecules in a gas, for instance.

### v2

@@ 5 Results and discussion
Figure 1: Affirmation frequency for responses to each question. The bars represent different samples: all (black), undergraduate (yellow), graduate (blue), Wikipedia notable (red), pre-1980 (grey), post-1980 (green), and post-2000 (magenta). The error bars reflect the effect of including values inferred for logical consistency.
Figure 2: Rejection frequency for responses to each question. The bars represent different samples: all (black), undergraduate (yellow), graduate (blue), Wikipedia notable (red), pre-1980 (grey), post-1980 (green), and post-2000 (magenta). The error bars reflect the effect of including values inferred for logical consistency.
We begin by discussing propositions related to the nature of the quantum state. Interestingly, we see that not all the books regard $\psi$ as ontologically complete. However, the support for this proposition is robust with assent frequency $0.85$ in the full sample, with all subsamples having assent frequency $\geq 0.8$. It is interesting that the lowest support for this occurs within the undergraduate and post-2000 samples, with outright rejection passing $0.1$ for post-1980 and post-2000 subsamples. Notably, there is no rejection in the pre-1980 sample, giving an indication that responses to this proposition are evolving. In terms of dissent, Phillips [32] and Moore [36] limit themselves to saying that $\psi$ can determine all the probabilities associated with a system but do not go on to say this is all the information allowed. Whereas, [21] openly suggests that interpretive difficulties might be overcome by finding a more general theory to supersede quantum mechanics.
We also see strong support (a minimum frequency of $0.94$) for the notion that a quantum state describes a single instance (as opposed to an ensemble). Notably, this achieves $0.96$ assent in the full sample and the undergraduate works are near unanimous on this count. In state collapse we find an interesting case. Here, both the assent and rejection frequencies increase as we move from pre-1980 to more modern subsamples. This seems to indicate the point is becoming more polarising over time. The collapse proposition still has subsample assent frequencies above $0.75$ while achieving $0.79$ in the full sample.
From these results we can also see that an orthodox view requires the state to contain the entire ontological situation and apply to individual particles. This, in turn, requires us to believe that superposition states can be occupied by individual systems. Of the three propositions, only affirmation of collapse reduces slightly if we consider sources after 1980. There are tentative signs of evolution in the consensus here. Particularly, the rate of rejection for collapse and ontological completeness both increase substantially when the post-1980 (and post-2000) subsample is compared to the full set. However, the assent frequency for collapse also increases from pre to post-1980, perhaps a marker of increased controversy. It is also noteworthy that, in a sample of $42$ works, there are only three dissenters to the notion that quantum states collapse: Weinberg [21], Merzbacher [51], as well as Binney & Skinner [22]. Additionally, the topic is not mentioned by Peres, Mandl, Müller-Kirsten, and Schiff [38, 39, 29, 57]. Only Peres [38] dissents from the notion that quantum states describe individual instances. Interestingly, we note that Home [6] argues that collapse is not essential to characterise the orthodoxy. However, this is strongly disputed by the data gathered here. Overall, we see that the orthodoxy has it that the quantum state is ontologically complete, refers to individual instances of systems, and collapses on measurement.
Now we turn to propositions on the nature of particles. A strong degree of consensus is clear around the notion that individual particles display wave-like behaviour with assent frequency $0.88$ in the full sample, while maintaining subsample frequencies $\geq 0.78$. The only dissenters here being Weinberg [21], Binney & Skinner [22], and Sakurai [42]. The topic is also not discussed in von Neumann [55] (perhaps unsurprising given the mathematical focus) and Griffiths [37]. Interestingly, the lowered assent in the post-1980 subsample seems to stem from more recent graduate texts rejecting the proposition. There is also notable evolution, in the sense that there are no rejections pre-1980, with these increasing in frequency as we move to post-1980 and post-2000 subsamples.
Reasonable support also exists for the notion of wave-particle duality with assent $0.78$ in the full sample and subsample frequency $\geq 0.71$. Here only von Neumann [55], Binney & Skinner [22], and Tong [16] outright deny this proposition. It is also not mentioned by Weinberg [21], Mandl [39], Townsend [40], Davydov [49] and Sakurai [42]. There is discernable evolution based on publication year, rejection is much more common in the post-2000 subsample than either post-1980 or pre-1980 cases.
For the proposition that individual particles being subject to the uncertainty principle, that is, one cannot know a given particle’s position and momentum (for example) simultaneously to arbitrary precision, we see more controversy. This is so because the full sample finds assent of $0.85$ and the pre-1980 subsample has nearly $1.0$. However, the post-1980 subsample only registers $0.77$, this then resurges to over $0.9$ in the post-2000 subset. This proposition receives the greatest rejection frequency in the post-1980 subsample. However, it is not rejected at all in the pre-1980 or post-2000 cases. The works between 1980 and 2000 that cause this are Griffiths [37], Peres [38], McIntyre [24], Moore [36], and Sakurai [42]. This shows some evidence of a change in the consensus over time, but it appears to have been a fluctuation, returning towards the pre-1980 consensus after two decades. Home [6] does not include this proposition in their conception of the orthodoxy, likely as it is a flawed proposition. As argued in [9], the uncertainty principle could not function in the suggested manner. If the measurements from individual particles had the uncertainty principle’s scatter, the final ensemble distribution would have a larger uncertainty. Thus, preventing the verification of a lower bound. Despite scepticism on this point, it is clearly wrong to exclude this from the orthodoxy, as the idea is uniformly affirmed by older works and is resurgent in recent texts. It is possible that this could be connected to the rise of interpretations of QM that use “information” as an explanatory tool. A common formulation of the uncertainty principle in terms of “allowed information” then becomes very tempting.
It is notable that affirmation of complementarity is generally weak, even the full sample only finds an agreement frequency of $0.52$. This is not due to denial (no one actually rejects it explicitly) but simply never being mentioned in many later texts. This accounts for the sustained drop in assent frequency from pre-1980 to post-1980 and post-2000 samples. Regardless, it cannot have a place in any robust orthodoxy. We also see that the issue of whether particles possess trajectories has become contentious, with frequency $0.51$ in the full sample. This becomes even more contentious in the post-1980 and Wikipedia samples with $0.44$ and $0.39$ respectively. It initially enjoyed an assent frequency $\sim 0.6$ pre-1980 but has dropped to around $\sim 0.4$ in more recent works. The idea is never rejected pre-1980, but rejection appears afterwards and is largely maintained in more recent works. Asserting there are no trajectories still finds agreement in modern works like Kok [20], Miller [27], Jain [28], Basdevant [31], Konishi & Paffuti [26], Cohen-Tannoudji [45], Shankar [43], Greiner [41], and Phillips [32]. There is also partial agreement in Moore [36] and Longair [23]. However, the mode response is 0 in the undergraduate, ‘Wikipedia notable’, and post-1980 samples. So there is some indication that this proposition is increasingly marginalised by the orthodoxy. It might also follow that those who affirm the “single particle uncertainty” proposition should, for consistency, agree that particles have no trajectory (at least for resolutions finer than that allowed by uncertainty). However, we have tried to avoid forcing logical consistency on authors where there is some ambiguity or nuance.
@@
Next, we turn to propositions around the question of why quantum mechanics predicts only statistical results. There is clear consensus, regardless of whether we confine our attention to modern works, that measurement disturbs quantum systems with assent $0.91$ in the full sample, with a minimum of $0.81$ in the graduate case. Notably, the pre-1980 and post-2000 samples have very similar assent frequencies. Indeed, only Peres [38] actively dissents. Their dissent is justified by the possibility of non-interactive measurement in QM [59].
@@
There is also support for the idea that measurement itself actually creates the observed values with frequency $0.76$ (without inferred values this drops to $0.59$) in the full sample while retaining 0.730.73($0.58$) in the post-1980 subsample. Interestingly, the assent frequency drops further in the post-2000 case, indicating that there is some evolution on this proposition. However, it is never rejected in any sample. It should be noted that the affirmation is only stated openly around 6060% of the time, despite being a plausible corollary to “measurement disturbs quantum systems” and that there are no active dissenters, not even [38].Only three authors do not provide grounds for inferring agreement: Sakurai [42], Mandl [39], and Weinberg [21]. The common case for inference is as follows: $\psi$ applies to individual systems, hence superpositions are real. Measurement disturbs the system and changes a superposition to an eigenstate, so it is either changing an unknown underlying value or creating a definite one where there was none. In either case the pre-existing value is totally superfluous if the collapse process is irreversible and indeterminate. However, the statistics here are significantly affected by the reliability of the inferences being made.
The notion that probabilities in QM result from the uncertainty principle has always been a minority view it appears, as most authors do not mention anything like it, achieving a maximum assent of $0.4$ in the pre-1980 sample. However, it is not rejected until after 1980 with the rejection rate dropping substantially in the post-2000 works. This coincides with an increase in assent post-2000, where Tong [16], Tamvakis [19], Longair [23], Konishi & Paffuti [26], and Müller-Kirsten [29] all support the position. In the same sample, only Binney & Skinner [22] reject it, so the mode response is still to leave it out.
There is still some support for the notion that the uncertainty principle is a consequence of measurement disturbance, with assent frequency $0.69$. It falls from $0.8$ pre-1980 to around $0.6$ post-1980 and 2000. Thus, providing clear evidence for consensus evolution on this point. Interestingly, the sample with most affirmation for this last proposition is the “Wikipedia notable” one, with $0.79$ assent. The only outright rejections here are Rae [15] and Basdevant [31], who both note the uncertainties can happily be measured in separate runs of an experiment, so “disturbance” need not occur.
@@
Thus, for the orthodoxy, we do not reveal the values of studied quantum systems but force them to assume definite values (or alter them) via some uncontrolled disturbance attached to the abstract notion of measurement. The “uncontrolled” aspect of this is vital to square the statistical nature of QM with the proposition that quantum states apply to individual instances of systems.
Next, we can ask whether there are correlations between proposition responses. To do this, we determine the frequency with which answers for different pairs of propositions match. We present the results in Table 3 for the full and post-1980 samples of textbooks. We choose only to display those for which the assent frequency exceeds $0.6$ to determine if there is a strong correlation between orthodox positions. Importantly, the strongest correlation occurs for the two pillars of the orthodoxy: “measurement disturbs systems” and “non-ensemble $\psi$”, indicating that they are correlated 9595% of the time. Over all, it is evident that relatively strong pair-wise correlations exist between all the propositions with assent frequency above $0.6$. They all correlate strongly to one (or both) of the two pillar propositions as well. This demonstrates that selecting an orthodoxy on this basis yields a consistently affirmed set of propositions.
Proposition 1 Proposition 2 Full sample Post-1980
Collapse Non-ensemble $\psi$ 0.81(0.79) 0.81(0.81)
Collapse Ontologically complete 0.76(0.74) 0.74(0.74)
Collapse Wave-particle duality 0.69(0.67) 0.74(0.74)
Collapse Single particle uncertainty 0.71(0.69) 0.67(0.67)
Collapse Single particle wave-like 0.79(0.76) 0.81(0.81)
Collapse Measurement creates values 0.81(0.52) 0.81(0.56)
Collapse Measurement disturbs system 0.81(0.76) 0.81(0.78)
Non-ensemble $\psi$ Ontologically complete 0.83(0.81) 0.81(0.78)
Non-ensemble $\psi$ Wave-particle duality 0.74(0.74) 0.70(0.70)
Non-ensemble $\psi$ Single particle uncertainty 0.90(0.88) 0.85(0.81)
Non-ensemble $\psi$ Single particle wave-like 0.86(0.83) 0.81(0.78)
Non-ensemble $\psi$ Measurement creates values 0.90(0.57) 0.85(0.56)
Non-ensemble $\psi$ Measurement disturbs system 0.95(0.90) 0.96(0.89)
Ontologically complete Wave-particle duality 0.64(0.64) 0.63(0.63)
Ontologically complete Single particle uncertainty 0.79(0.76) 0.74(0.70)
Ontologically complete Single particle wave-like 0.79(0.79) 0.78(0.78)
Ontologically complete Measurement creates values 0.81(0.52) 0.78(0.48)
Ontologically complete Measurement disturbs system 0.81(0.79) 0.81(0.78)
Wave-particle duality Single particle uncertainty 0.67(0.67) 0.59(0.59)
Wave-particle duality Single particle wave-like 0.76(0.76) 0.74(0.74)
Wave-particle duality Measurement creates values 0.83(0.55) 0.85(0.63)
Wave-particle duality Measurement disturbs system 0.74(0.71) 0.74(0.70)
Single particle uncertainty Single particle wave-like 0.83(0.81) 0.78(0.74)
Single particle uncertainty Measurement creates values 0.83(0.50) 0.74(0.44)
Single particle uncertainty Measurement disturbs system 0.86(0.83) 0.81(0.78)
Single particle wave-like Measurement creates values 0.86(0.55) 0.81(0.56)
Single particle wave-like Measurement disturbs system 0.83(0.81) 0.81(0.78)
Measurement creates values Measurement disturbs system 0.90(0.55) 0.89(0.56)
Table 3: The frequency with which answers for two propositions match in the full and post-1980 samples. The values in brackets reflect the case without inference of responses.
@@
where the sum runs over the propositions, $w_{i}$ is the average response value, and $a_{i}$ is that of the reference work in question. Note that we exclude propositions that do not have a full-sample assent frequency >0.6>0.6. As illustrated above, this selection criterion results in set of well-correlated propositions.
The average orthodoxity is 7474% with standard deviation 2323% indicating a robust degree of agreement on average. This includes inferred values, but, this effect is small, going to 7171% and 2424% without them. Note that these averages are robust to increasing the frequency inclusion threshold to $0.7$ as well. For convenience, we order Fig. 3 by year of first publication. This allows us to determine if there are any trends in the orthodoxy. Early works adhere very closely to the full-sample orthodoxy, with those before 1980 having an average orthodoxity of 8383% with standard deviation 1414%. The period around the early 90’s shows a fluctuation of far lower agreement with the orthodoxy. This results in post 1980 works having an average orthodoxity of 6868% and standard deviation 2525%. The substantial nature of this change from the earlier consensus is most vividly illustrated by the 50% increase in the post-1980 standard deviation, despite a 7373% increase in sample size. The 90’s dip in the orthodoxity is constituted by the works: Peres [38], Moore [36], Griffiths [37], Mandl [39], Rae [15], and Sakurai [42]. Of these, Mandl is here mainly due to omission, rather than contesting orthodox positions. If we replace Mandl with Weinberg [21], we form a set of all the authors who deny more than one proposition. However, there is no consistent pattern in these denials.
Overall, there appears to be a post-1980 divergence from the earlier orthodoxy. This being said, it resembles a fragmentation, rather than a new consensus emerging, due to the lack of consistent patterns in the denial of previously orthodox positions. However, it must be admitted that many modern works adhere quite closely to the interpretation put forward by their predecessors. This suggests the possibility that our present epoch is the start of a collapse in confidence in the orthodoxy. A survey of working physicists has recently demonstrated that the community is fragmented on the interpretation of quantum mechanics [2], notably more so than a previous survey in 2016 [1]. This could well signal an increase in divergence from the previous quantum orthodoxy in future textbooks. Somewhat complicating such an idea is the tendency of more orthodox post-1980 works to cite Bell’s theorem as a vindication of many aspects of the orthodoxy, perhaps accounting for the apparent resurgence in orthodoxity after the late 1990’s. The timeline notably aligns with the emergence of “Copenhagen+” interpretations like quantum Bayesianism (see [60]).
@@
For additional works that were not included in the sample we find the following orthodoxity measures: Ballentine [14] 0.88-0.88, Home [6] 0.42, Hannabuss [17] 0.45, and Nielsen & Chuang [18] 0.42. The last of these, Nielsen & Chuang [18], only shows a lower orthodoxity by virtue of never mentioning the topic of waves. Their strong endorsement of Peres [38] and Sakurai [42] might indicate little sympathy with this aspect of the orthodoxy, but the clarity of this is marred by also endorsing Cohen-Tannoudji [45]. Interestingly, Home’s characterisation of the orthodoxy [6] appears incomplete at $0.42$. Additionally, this is only when one concentrates on the aspects they attributed to the non-ensemble case. Home considers ensemble views as part of the orthodoxy on the basis that they view the formalism as being complete (there are no supplemental entities needed to explain what is going on). This principle as the central pillar of the orthodoxy is apriori dubious. This is because the notion of collapse is a supplement to the basic mathematical formalism [61, 14, 21], in the form of an additional physical/quasi-physical process with attendant projection postulate. Since Copenhagen-like interpretations hinge so heavily on collapse, their popularity implies that Home’s primary condition of orthodoxy becomes more doubtful. Interestingly, Ballentine [14], who would express an orthodox view according to Home, achieves an orthodoxity of 0.88-0.88 here, dissenting from (or not addressing) every proposition considered. This is because Home’s rationale for including ensemble views in the orthodoxy may not be well justified. Home argues that, to enjoy a complete formalism, the ensemble must be supplemented by hidden variables that explain the assignment of definite values to individual realisations. However, an ensemble view only requires statistical completeness, $\psi$ can answer all statistical questions, not true ontological completeness. In this regard Home forces ensemble views to be able to determine the behaviour of single instances, in contradiction with their basic premise. In addition, no such definite assignment is required of other statistical theories in science (which usually exist as a limit of some other, more detailed theory). We do not require statistical mechanics to explain how velocities are assigned to the molecules in a gas, for instance.